# Optimized Many-Hypercube Codes toward Lower Logical Error Rates and Earlier Realization

### v1

@@ VI Summary and outlook
Toward lower logical error rates and earlier experimental realization of MHC codes, we have investigated the MHC codes with $[[6,4,2]]$ and/or $[[4,2,2]]$ quantum error-detecting codes. Consequently, we have found that $D_{6,4,4}$ and $D_{6,6,4,4}$ achieve the lowest decoding error probability among the level-3 and level-4 MHC codes, respectively. Focusing on level 3, we have also developed efficient fault-tolerant encoders achieving about 60% overhead reduction compared with the original encoders. The proposed encoders also improve the logical-CNOT performance a little. $D_{6,4,4}$ has also showed the highest performance for logical CNOT gates. Thus, we conclude that $D_{6,4,4}$ is the best of the level-3 MHC codes. This code can achieve not only lower logical error rates but also the smaller number of necessary physical qubits compared to the original $D_{6,6,6}$. Therefore, $D_{6,4,4}$ will become a good target toward experimental realization of FTQC with a high-rate quantum code.
In this work, we have evaluated only logical CNOT gates and left the evaluation of other logical gates for future work. Also, we have assumed no memory errors in the present circuit-level noise model. Toward experimental realization, e.g., with trapped ions or neutral atoms, we should optimize the gate sequence to minimize the moving distance of qubits. Such optimization is also left for future work.

### v2

@@ VI Summary and outlook
Toward lower logical error rates and earlier experimental realization of MHC codes, we have investigated the MHC codes with $[[6,4,2]]$ and/or $[[4,2,2]]$ quantum error-detecting codes. Consequently, we have found that $D_{6,4,4}$ and $D_{6,6,4,4}$ achieve the lowest decoding error probability among the level-3 and level-4 MHC codes, respectively. Focusing on level 3, we have also developed efficient fault-tolerant encoders realizing about 60% overhead reduction compared with the original encoders. The proposed encoders also improve the logical-CNOT performance a little. It have turned out that $D_{6,4,4}$ shows the highest performance for logical CNOT gates. Thus, we conclude that $D_{6,4,4}$ is the best of the level-3 MHC codes. This code can achieve not only lower logical error rates but also the smaller number of necessary physical qubits, compared to the original $D_{6,6,6}$. Therefore, $D_{6,4,4}$ will become a good target toward experimental realization of FTQC with a high-rate quantum code.
In this work, we have evaluated only logical CNOT gates and left the evaluation of other logical gates for future work. Also, we have assumed no memory errors in the present circuit-level noise model. Toward experimental realization, e.g., with trapped ions or neutral atoms, we should optimize the gate sequence to minimize the moving distances of qubits. Such optimization is also left for future work.